### v1

# SensorChat: Answering Qualitative and Quantitative Questions during Long-Term Multimodal Sensor Interactions

Abstract.
Natural language interaction with sensing systems is crucial for enabling all users to comprehend sensor data and its impact on their everyday lives. However, existing systems, which typically operate in a Question Answering (QA) manner, are significantly limited in terms of the duration and complexity of sensor data they can handle.
In this work, we introduce SensorChat, the first end-to-end QA system designed for long-term sensor monitoring with multimodal and high-dimensional data including time series. SensorChat effectively answers both qualitative (requiring high-level reasoning) and quantitative (requiring accurate responses derived from sensor data) questions in real-world scenarios. To achieve this, SensorChat uses an innovative three-stage pipeline that includes question decomposition, sensor data query, and answer assembly. The first and third stages leverage Large Language Models (LLMs) for intuitive human interactions and to guide the sensor data query process. Unlike existing multimodal LLMs, SensorChat incorporates an explicit query stage to precisely extract factual information from long-duration sensor data. We implement SensorChat and demonstrate its capability for real-time interactions on a cloud server while also being able to run entirely on edge platforms after quantization. Comprehensive QA evaluations show that SensorChat achieves up to 26% higher answer accuracy than state-of-the-art systems on quantitative questions. Additionally, a user study with eight volunteers highlights SensorChat’s effectiveness in handling qualitative and open-ended questions.
Question Answering, Multimodal Sensors, Large Language Models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Computer systems organization Embedded systemsccs: Computing methodologies Machine learningccs: Computing methodologies Natural language processing
1. Introduction
In recent years, the number of Internet of Things (IoT) devices in daily life has grown rapidly, including smartphones, smart home accessories, and smart city infrastructures (Choi et al., 2021; Khang et al., 2023). These devices generate a massive amount of multimodal sensor data, with total production expected to reach 80 ZB by 2025 (International Data Corporation (2021), IDC). Traditional research has focused on developing classification models for tasks such as human activity recognition (Ouyang et al., 2022, 2023; Xu et al., 2023b, a; Cao et al., 2024; Weng et al., 2024). In contrast, latest studies have recognized the importance of natural language interactions, especially for making sensor data more accessible and useful to elderly people or patients with Alzheimer’s (Xing et al., 2021; Nie et al., 2022b; Yang et al., 2024b; Arakawa et al., 2024; Ji et al., 2024; Chen et al., 2024). These systems typically operate in a Question Answering (QA) framework: users ask questions, the sensor system processes both the query and the underlying sensor data, and finally delivers an adequate answer. The questions can be qualitative, requiring high-level reasoning about a user’s lifestyle or well-being (e.g., “How is my work-life balance?”), or quantitative, demanding quantitatively precise answers based on sensor data (e.g., “How long did I walk?”). QA offers a natural and flexible way for users to interact with and gain insights from sensors, going beyond simple classification models.
(a) Example QA in Sensor2Text (Chen et al., 2024).
(b) Example QA in DrHouse (Yang et al., 2024b).
(c) Example quantitative (left) and qualitative (right) QA in SensorChat (our system).
Figure 1. A visual comparison of QA in state-of-the-art sensor systems and SensorChat. SensorChat excels in processing long-duration sensor data and generating quantitatively accurate answers and insightful qualitative analyses of well-being.
However, existing QA systems for sensors face significant limitations, particularly with short time windows or low-dimensional sensors. Most current systems can only process data over short periods, such as one minute in DeepSQA (Xing et al., 2021) or 16 seconds in Sensor2Text (Chen et al., 2024), as illustrated in Fig. 1(a). While useful for both qualitative and quantitative activity analysis, such short time frames offer limited value for gaining deeper insights into users’ fitness and overall well-being. Other QA systems focus on longer-term clinical diagnosis or health-related queries, but leverage low-dimensional physiological data like average heart rate or daily step counts (Englhardt et al., 2024; Kim et al., 2024; Yang et al., 2024b), as shown in Fig. 1(b). While effective for certain use cases, they cannot handle high-dimensional time series data, such as IMU signals, and fail to provide accurate answers about fine-grained activities. In practice, reasoning based on fine-grained physical activities, such as daily time spent exercising versus sitting, is critical for assessing a user’s health status (of Health et al., 2021). A significant gap remains in developing a QA system capable of answering both qualitative and quantitative questions using (1) long-duration and (2) high-dimensional multimodal sensor data.
In this paper, we introduce SensorChat, the first end-to-end system designed for answering both qualitative and quantitative queries in long-term sensor monitoring, using multimodal and high-dimensional time series sensors. We consider the commonly used mobile devices for sensor data sources, i.e., smartphones and smartwatches. In contrast to prior QA systems (Xing et al., 2021; Nie et al., 2022b; Chen et al., 2024; Arakawa et al., 2024; Yang et al., 2024b), SensorChat is capable of processing long-duration sensor data spanning weeks or even months. For example, Fig. 1(c) shows a user’s activities over a three-day period, with the x-axis representing wall-clock time from 00:00 AM to 11:59 PM, and the y-axis denoting separate days. Different activity labels are depicted by color-coded bars. When asked a quantitative query like, ”What day did I spend the longest eating last week?” SensorChat provides a precise answer based on a detailed search of the sensor database, such as, ”You spent the most time eating on Saturday.” Furthermore, leveraging its detailed analysis of daily routines, SensorChat can also properly address open-ended qualitative queries related to overall health and lifestyle, such as, ”How would you rate my lifestyle?” To the best of the authors’ knowledge, SensorChat is the first QA system capable of processing such long-duration sensor data to provide both precise activity-based responses and insightful evaluations of lifestyle trends.
To achieve the goal, SensorChat introduces a novel three-stage design including question decomposition, sensor data query and answer assembly. The question decomposition and answer assembly stages utilize the power of Large Language Models (LLMs) to correctly understand arbitrary questions and provide accurate answers. SensorChat further enhances the LLMs in these stages via in-context learning and finetuning techniques to adapt LLMs for various QA scenarios. The intermediate sensor data query stage is pivotal in encoding raw time series data from multimodal sensors into a pre-stored embedding database. During offline pretraining, SensorChat uses a novel contrastive sensor-text pretraining loss to align sensor and text encoders. In the online stage, SensorChat accurately and exhaustively retrieves all relevant sensor embeddings by performing a similarity search against text embeddings generated from question decomposition. These queried embeddings are then processed by a set of carefully designed summarization functions to generate the sensor context in text. Finally, in the answer assembly stage, the sensor context is integrated with the original question using another LLM, to generate the ultimate answer. The explicit query mechanism and the fusion of textual sensor information with user questions distinguish SensorChat from emerging multimodal LLMs, which often rely on the ”black-box” integration of natural language and sensory data (Moon et al., 2023b; Zhang et al., 2023; Liu et al., 2024a; Han et al., 2024). As a result, SensorChat excels in answering long-term qualitative and quantitative questions accurately, where state-of-the-art QA systems fall short (Xing et al., 2021; Nie et al., 2022b; Chen et al., 2024; Arakawa et al., 2024; Yang et al., 2024b).
We implement SensorChat in two system variants to accommodate various user needs: SensorChatC (“C” stands for “Cloud”) for deployment on a cloud server and SensorChatE (“E” stands for “Edge”) for deployment on an edge device. SensorChatC leverages OpenAI’s GPT APIs (Ye et al., 2023; Achiam et al., 2023) for better interpretation of user queries, enabling real-time QA interactions on a server equipped with an NVIDIA A100 GPU (a10, 2025). On the edge side, SensorChatE operates entirely on an edge device, preserving user privacy albeit at the cost of increased latency. We test SensorChatE on an NVIDIA Jetson Orin NX module (jet, 2025).
We conduct comprehensive evaluations of SensorChat on a state-of-the-art QA dataset, mainly covering the quantitative queries. The dataset is carefully chosen to include long-term multimodal sensor data and practical QA samples aligned with users’ interests. Additionally, we recruited eight volunteers for a real-world user study, approved by the Institutional Review Board (IRB). Participants were instructed to carry a smartphone while following their daily routines for one to three days. At the end of the sensor data collection period, they were encouraged to interact with SensorChat, ask any questions they had, and provide their feedback.
In summary, the major contributions of our work are:
  • We propose SensorChat, the first end-to-end system for answering qualitative and quantitative questions in long-term monitoring using high-dimensional time series sensors. SensorChat is designed with a novel three stage pipeline, including question decomposition, sensor data query and answer assembly.
  • We introduce a novel contrastive sensor-text pretraining loss for training the sensor and text encoders offline, enabling accurate sensor embedding queries given arbitrary partial text during the online sensor data query.
  • We enhance the LLMs in SensorChat through in-context learning and fine-tuning, incorporating carefully designed prompts to improve the quality of question decomposition and answer assembly.
  • Comprehensive evaluations show that SensorChat improves answer accuracy by up to 26% on quantitative questions, compared to the best-performing state-of-the-art QA systems. A real-world user study confirms SensorChat ’s ability to answer qualitative and open-ended questions to users’ satisfaction. In terms of latency, SensorChatC enables real-time user interactions from a cloud server, while SensorChatE operates fully on local edge devices with higher latency.
2. Related Work
Question Answering using Sensor Data The QA problem has been extensively studied across various domains, including text (Rogers et al., 2023), visual (Schwenk et al., 2022), medical (Pal et al., 2022), and remote sensing (Hu et al., 2023). In the sensor domain, early works AI Therapist (Nie et al., 2022a) and its successor CaiTI (Nie et al., 2024) utilized smart home devices, such as Amazon Echo, to engage in conversations with users and assess mental well-being. DeepSQA (Xing et al., 2021) was the first to benchmark time-series sensor-based QA for human activity recognition. It introduced SQA-GEN, an automated QA generation tool that gathers 1-minute sensor readings and generates valid Q&A pairs by exhaustively searching six pre-defined question templates. They mainly focused on quantitative questions including time query, counting and action compare. The authors also evaluated traditional neural network models, including CNNs and LSTMs, finding that the ConvLSTM network with Compositional Attention achieved the highest QA accuracy.
Recent contributions have pioneered the integration of LLMs for generating more intuitive answers and better interpreting sensor data. Englhardt et al. (Englhardt et al., 2024), Health-LLM (Kim et al., 2024), and DrHouse (Yang et al., 2024b) converted physiological data from wearable devices, such as heart rates and daily step counts, into text prompts for LLMs, enabling more sophisticated medical and healthcare diagnoses. The latest Sensor2Text (Chen et al., 2024) and PrISM-Q&A (Arakawa et al., 2024) explored natural language interactions between users and wearable devices to understand and support daily activities using sensor data, such as advising on ”What should I do next with this?” Both approaches utilized LLMs as their backbone and fed sensor embeddings into the models.
While promising, existing systems exhibit significant limitations in handling long duration and complex sensor data required for accurate answers. Table 1 highlights the key differences between SensorChat and existing QA systems. In summary, SensorChat greatly enhances capabilities of existing systems by (1) answering questions based on long-duration sensor data spanning weeks or months, compared to the short windows of seconds or minutes in previous systems (Xing et al., 2021; Chen et al., 2024; Arakawa et al., 2024), and (2) encoding high-dimensional time-series data to extract fine-grained activity details for reasoning, unlike prior systems that are limited to low-dimensional sensor data (Englhardt et al., 2024; Kim et al., 2024; Yang et al., 2024b).
LLMs for Multimodal Reasoning Recent works investigated multimodal LLMs that transform other data modalities into a sequence of tokens for LLM integration (Zhang et al., 2023; Lin et al., 2024b). IMU2CLIP (Moon et al., 2023b) and TENT (Zhou et al., 2023) employed contrastive pretraining to align text with various timeseries sensor signals. Recent designs like AnyMAL (Moon et al., 2023a) and OneLLM (Han et al., 2024) proposed fine-tuning multimodal LLMs to process up to eight different modalities, including IMU time series, for reasoning tasks. However, direct integration of time series sensor data to LLMs is constrained by LLM’s inherent weakness in handling long-context inputs (Li et al., 2024; Gu and Dao, 2023), making them unsuitable for processing long-duration sensor signals. In contrast, SensorChat overcomes this limitation by incorporating a dedicated sensor data query stage, enabling it to handle long-term queries effectively.
Table 1. Comparing SensorChat and existing QA systems.
Existing QA systems for sensor data Long-duration High-dimensional
sensor data time series sensors
DeepSQA (Xing et al., 2021), Sensor2Text (Chen et al., 2024), PrISM-Q&A (Arakawa et al., 2024) $\times$
Englhardt et al. (Englhardt et al., 2024), Health-LLM (Kim et al., 2024), DrHouse (Yang et al., 2024b) $\times$
SensorChat (this work)
Mobile Systems for Daily Life Monitoring Researchers have explored a variety of sensing technologies for monitoring human lives, including built-in sensors on smart phones (Zhang et al., 2020), cameras (Radu and Henne, 2019), Wi-Fi signals (Yang et al., 2024a) and mmWave radar (Weng et al., 2024). Although these efforts have resulted in numerous open-source datasets (Reyes-Ortiz et al., 2012; Banos et al., 2014; Roggen et al., 2012; Vaizman et al., 2017) and powerful machine learning models (Ma et al., 2019; Xu et al., 2021; Deldari et al., 2022; Ouyang et al., 2022, 2023; Xu et al., 2023b), they fail to handle queries in natural language, which are more creative and open-ended. In this paper, we design SensorChat to facilitate long-term QA interactions based on time series sensor data. The methodology of SensorChat can be further integrated with other sensing modalities and applications.
3. Background and Motivation
In this section, we formally define the problem we address and present a motivating study highlighting the limitations of state-of-the-art QA techniques.
3.1. Problem Statement
We target real-world QA scenarios in which users carry mobile sensing devices during their daily routines, with minimal attention to the sensing process. In this context, sensor data is collected continuously over an extended period in natural, uncontrolled environments. During this process, users may pose arbitrary, personalized questions about the sensor data, ranging from inquiries about a single day to those spanning several weeks, and expect precise answers derived from the collected data.
Formally, the problem is formulated as generating a proper answer given the question and sensory inputs, denoted as $\mathbf{x}$. The questions can be arbitrary, including qualitative and quantitative queries, while the answers are expected to accurately address users’ concerns. The key distinction between our problem formulation and existing QA approaches lies in the format of $\mathbf{x}$, which consists of long-duration high-dimensional data collected up to the moment the question is asked. As shown in Table 1, this makes our problem significantly more challenging.
3.2. Limitations of SOTA Methods
With a concrete problem setup, we conduct a motivating study to understand the limitations of state-of-the-art systems on this problem. We then discuss the lessons learned which inspire the design of SensorChat.
Dataset Setup We choose the SensorQA dataset [^1] https://github.com/benjamin-reichman/SensorQA (Reichman et al., 2025) to simulate long-term real-world QA interactions between humans and sensing devices. To our knowledge, SensorQA is the first and only existing QA dataset that matches our target scenario, i.e., in-the-wild, long-term, time series sensor data collection with practical questions that align with users’ interests.
Specifically, SensorQA (Reichman et al., 2025) used the ExtraSensory dataset (Vaizman et al., 2017, 2018b) as its sensor data source, which contains raw sensor measurements from 60 subjects and has more than 300K minutes of data, annotated with 51 activity or context labels after cleaning. The data were collected using easily accessible smartphone and smartwatch sensors, including IMUs, audio (MFCC features), and phone state sensors (compass, location, Wi-Fi, light, etc.). Unlike other human activity recognition datasets (Reyes-Ortiz et al., 2012; Banos et al., 2014; Roggen et al., 2012; Ouyang et al., 2023), ExtraSensory is notable for its natural and uncontrolled collection environment, with no restrictions on device placement.
The QA data in SensorQA (Reichman et al., 2025) were generated by real human workers as if they were interacting with a smart sensor device about their daily lives. As a result, SensorQA (Reichman et al., 2025) comprises diverse and practical questions that provide valuable insights for users. SensorQA covers a wide range of topics from normal activities to work patterns to work-life balance, and question types such as time queries, counting, and activity recognition. In SensorQA (Reichman et al., 2025), the authors identified six question categories and seven answer categories. Table 2 details the categories along with example Q&As and the corresponding number of instances. While SensorQA exhibits a skewed distribution favoring quantitative and time-related queries, this trend underscores users’ strong interest in gaining temporal insights into their lives, highlighting the need for a QA system to address quantitative queries.
Question Example Questions # of
Categories Questions
Time Compare Did I spend more time sitting or standing? 1,432
Day Query On which day did I spend the most time at home? 1,277
Time Query How long was I in class and at school? 1,119
Counting How often did I groom? 725
Existence Did I have a meeting on Wednesday? 668
Action Query What did I do after I left home on Tuesday? 428
(a) Question categories.
Answer Example Shortened # of
Categories Answers Answers
Action Doing computer work 1,357
Day/Days Last Friday 1,242
Existence Yes/No 1,047
Time Length 40 Minutes 1,018
Location At school 792
Count Three times 401
Timestamp Around 11:00 am 310
(b) Answer categories.
Table 2. Q&A categories in the SensorQA dataset (Reichman et al., 2025). The short answers are presented for simplicity.
In total, SensorQA (Reichman et al., 2025) contains 5.6K QA pairs reflecting practical human interests in daily life monitoring, including about 3K questions on daily activities and 2.6K on longer durations spanning weeks or months. We randomly split the QA data in SensorQA (Reichman et al., 2025) into an 80/20 train-test set. During the experiments, the full sensor data history for each user and one question are fed to the baselines. The generated answer is compared with the ground-truth answer in SensorQA (Reichman et al., 2025).
State-of-the-art (SOTA) Baselines Setup In this motivating study, we select three state-of-the-art methods that have open-sourced code, each representing one mainstream category of approaches:
  • DeepSQA (Xing et al., 2021) represents the SOTA CNN-LSTM based model for fusing the question and time series sensor data. We specifically train and test the DeepSQA-CA model which uses Compositional Attention with CNN-LSTM architecture and achieves the highest accuracy as reported in the original paper. We use full-history IMU signals as sensor inputs, since the DeepSQA-CA model was designed to handle IMU time series data.
  • IMU2CLIP+GPT-4 (Moon et al., 2023b) is the SOTA GPT-based generative method. It uses IMU2CLIP (Moon et al., 2023b) to associate raw IMU signals with text and uses GPT-4 for reasoning and answer generation. For each timestamp, IMU2CLIP searches for a text description that best matches the IMU data. The complete narrative text from all timestamps is then fed into GPT-4 (Achiam et al., 2023) with the question to generate an answer. We train the IMU2CLIP model using the IMU sensor data in SensorQA (Reichman et al., 2025).
  • LLaMA-Adapter (Zhang et al., 2023) is the SOTA multimodal LLM model that integrates image or video inputs with the LLaMA backbone using a Transformer-based adapter. We use the activity Gantt charts as image inputs, representing the sensor information. We start with their pretrained weights on a LLaMA2-7B (Touvron et al., 2023a) backbone and then finetune it with LoRA (Hu et al., 2021) on SensorQA (Reichman et al., 2025).
(a) Example of single-day query.
(b) Example of multi-day query.
Figure 2. Ground-truth QA pairs in SensorQA (Reichman et al., 2025) and SOTA methods’ answers.
Lesson 1: LLMs are necessary for natural interactions, but not enough We first visualize two examples of the question, true answer and the response generated by the three SOTA methods in Fig. 2 (a) and (b). Similar as in Fig. 1, we visualize the user’s activity in Gantt charts. In Fig. 2 (a), we study a single-day query asking the activities after waking up on a specific day of Wednesday. While the answers are obvious to humans from the chart, this question requires multi-step reasoning: the algorithm first needs to figure out the end time of ”waking up”, and then queries the activity right after this timestamp. As a result, traditional methods such as DeepSQA cannot follow these non-trivial reasoning steps needed to process the full history of time series data. Additionally, DeepSQA can only generate short phrases because the CNN-LSTM is trained for classification, learning to select the most suitable answer from a limited set of candidates. The other two methods based on LLMs are able to produce complete sentences, showing LLM’s crucial role in natural interactions with humans.
However, LLMs do not always succeed. In Fig. 2 (a), LLaMA-Adapter ends up with an irrelevant answer of ”went to school”, likely a random guess rather than a result of multimodal fusion. Multimodal LLMs like LLaMA-Adapter may underperform without extensive high-quality data (e.g., over 1M samples) for finetuning. In this example, only IMU2CLIP+GPT-4 proves its ability to generate the correct answer of ”showering” and ”in the toilet”. This can be attributed to IMU2CLIP’s success in associating the IMU signals with the correct text, and GPT-4 generating the correct reasoning from the full narrative text. Nevertheless, IMU2CLIP+GPT-4 has a significant limitation due to GPT’s restricted prompt length. This leads to limited narrative text and a reduced sensor duration that can be processed. Specifically, IMU2CLIP+GPT-4 struggles with multi-day queries, as shown in Fig. 2 (b). The question asks for comparing the frequency of meals on multiple days, and then identifying the day with the fewest meals. All three methods fail in this challenging task. IMU2CLIP+GPT-4 in particular refuses to give any valid answer.
Lesson 2: SOTA methods struggle with accurate quantitative answers over long-duration sensor data Although we recognize the critical role of LLMs in user interactions, it is well known that LLMs struggle with quantitative analyzes and numerical reasoning, especially for long-context inputs (Li et al., 2024; Gu and Dao, 2023). We study the precision of answers generated by SOTA methods on SensorQA (Reichman et al., 2025). To measure accuracy, we match the key phrases from the ground-truth answers with those in the generated answers. Specifically, SensorQA offers a shortened version of each answer by distilling 1-2 key words from the full answers using GPT-3.5-Turbo (Ye et al., 2023). A generated answer is considered correct if it includes the ground-truth short answer; otherwise, it is marked as incorrect.
Figure 3. Answer accuracy of state-of-the-art methods and SensorChat by question category (left) and answer category (right). Categories are arranged from easier questions/answers on the left to more challenging questions/answers on the right.
Fig. 3 shows the accuracy results of SOTA methods by question category (left) and answer category (right). One key observation is that different Q&A types present varying levels of difficulty for the baselines. For example, existence and time comparison questions are easier to answer due to the limited number of possible responses (e.g., Yes/No for existence questions). Counting and time queries are more challenging as they require quantitative analysis. Additionally, different methods exhibit varying sensitivity to question and answer types. IMU2CLIP+GPT-4 completely fails in counting questions, as highlighted by the red arrow in Fig. 3 (left). This can be credited to GPT-4’s difficulty with counting in long narrative texts. Similarly, as pointed by the red arrow in the right figure of Fig. 3, all three baselines struggle to generate correct answers for timestamps, underscoring the weaknesses of existing methods in quantitative reasoning over long-duration sensor data. This weakness stems from the essence of existing designs, where long-duration raw sensor signals are directly fed into large models for fusion with text. Our key insight is that sensor-text fusion requires a more carefully designed approach to handle the diverse queries and long-duration multimodal sensor data encountered in real-life scenarios. To address this, we propose SensorChat, a novel system with a three-stage pipeline that includes a dedicated sensor data query stage. As shown in Fig. 3, SensorChat significantly improves answer accuracy across all categories, particularly for quantitative tasks such as day comparisons and time queries.
4. SensorChat Design
4.1. Overview of SensorChat
To address the challenge of accurate QA over long-duration sensor data, we design a novel three-stage pipeline in SensorChat, consisting of question decomposition, sensor data query, and answer assembly. Specifically, we use LLMs in question decomposition and answer assembly, recognizing their essential roles in correctly interpreting user queries and generating natural language answers. The intermediate sensor query stage is the key component of SensorChat. First, the query stage uses a pretrained sensor encoder to effectively encode high-dimensional and long-duration multimodal sensor timeseries into meaningful embeddings. Second, the query stage performs a similarity search in the embedding space to retrieve all sensor information relevant to the original question. This ensures precise extraction of sensor context, making a significant contribution to the accuracy of the final answer. To the best of the authors’ knowledge, SensorChat is the first system to incorporate an explicit sensor data query stage for accurately handling sensor-based language tasks in long-term monitoring.
The complete pipeline of SensorChat is illustrated in Fig. 4. For example, given a question like, “How long did I exercise last week in the morning?” (① in Fig. 4), SensorChat first decomposes the question and generates specific sensor data queries using LLMs. These queries include details such as the context of interest, date and time ranges, and the summarization function (③). In this case, the query might specify a context of “do exercise,” a time span of “last week,” a time of day of “morning,” and a summarization function of CalculateDuration. To enhance the precision of question decomposition, SensorChat uses solution templates (②) with carefully designed prompts. Next, the sensor query stage encodes the activity text “exercise” into the embedding space (④) and retrieves sensor embeddings that are sufficiently similar to the text embedding (⑤). These sensor embeddings are encoded offline from the full-history raw sensor data. Both the sensor and text encoders are pretrained offline to align their outputs in the same embedding space, ensuring accurate sensor data retrieval. Therefore, only text encoding and similarity-based query searches are performed online. Additional properties, such as the date (“last week”) and time of day (“morning”), are used to constrain the sensor query range. The final step of the sensor data query stage involves summarizing the relevant sensor embeddings into a textual context with a summarization function (⑥). For instance, with the CalculateDuration function, the summarization step calculates the total duration of the retrieved sensor embeddings. This results in a sensor context such as: “Among all days last week, you exercised for 35 minutes on Monday morning and 55 minutes on Thursday morning.” The summarization function is identified during question decomposition. Finally, the original question and the sensor context are fed into the answer assembly stage, where a fine-tuned LLM generates the final answer (⑦). With the precise sensor context, the model produces an accurate response, such as: “You exercised a total of 1 hour and 30 minutes last week.” (⑧).
Figure 4. The system diagram of SensorChat including three stages.
In the following lines, we describe the key design elements of SensorChat: the offline contrastive sensor-text encoder pretraining (Sec. 4.2) and the online three-stage pipeline (Sec. 4.3).
4.2. Contrastive Sensor-Text Pretraining for Partial Contexts
The sensor and text encoders are crucial for effective sensor-text fusion, serving as a ”bridge” between high-dimensional sensor timeseries and semantic text. A high-quality sensor encoder ensures that SensorChat can accurately and comprehensively retrieve relevant sensor embeddings during the sensor data query stage. Similarly, a robust text encoder is essential for handling the arbitrary contextual text generated during the question decomposition stage.
However, achieving this level of alignment poses significant challenges due to the lack of suitable techniques. Existing pretraining methods, such as CLIP and its variants (Radford et al., 2021; Moon et al., 2023b), are effective for pretraining encoders with paired inputs, i.e., one piece of sensor data and one sentence. Unfortunately, encoders trained in this manner struggle to accurately identify similar sensor embeddings when provided with partial or arbitrary context instead of complete sentences. For instance, using CLIP, a piece of sensor data might align well with the full sentence ”The person is sitting and working on computers at school.” However, when a user is specifically interested in a partial context like ”working on computers,” the encoded text embedding may not closely match the original sensor embeddings, leading to reduced accuracy in sensor data query as we show in Sec. 6.6. To address this challenge, we introduce a novel contrastive sensor-text pretraining loss for partial contexts.
[…]
Figure 5. Visualization of the contrastive sensor-text pretraining in SensorChat.
We introduce a new pretraining loss to align sensor embeddings $\mathbf{z}^{s}_{t}$ and partial text embeddings $\mathbf{z}^{w}_{t}$. Different from CLIP (Radford et al., 2021), our loss function enables alignment between sensor data and all partial phrases, defined as follows:
(1) $\mathcal{L}=\sum_{t}{\frac{-1}{|w_{t}|}\sum_{w\in w_{t}}\log\frac{\exp(\mathbf{z}^{s}_{t}\cdot\mathbf{z}^{w}_{t}/\tau)}{\sum_{a\in A(w)}\exp(\mathbf{z}^{s}_{t}\cdot\mathbf{z}^{a}/\tau)}},$
[…]
Table 3. List of important notations.
Symbol Meaning Symbol Meaning
$\mathbf{x}_{t}$ Multimodal sensor data sampled at $t$ $w_{t}$ Partial context labels at $t$
$d$ Number of different sensors $T$ Total duration of sensor data
$\theta$ Sensor encoder $\phi$ Text encoder
$\mathbf{z}^{s}_{t}$ Sensor embedding of $\mathbf{x}_{t}$ $\mathbf{z}^{w}_{t}$ Text embedding of $w_{t}$
$A(w)$ Set of negative contexts of $w$ $\tau$ Temperature scalar
$f$ Trained similarity function between embeddings $h$ Threshold in query search
4.3. Three Stages in SensorChat
In this section, we explain the detailed designs in each stage of SensorChat, including question decomposition, sensor data query and answer assembly. All stages need to work accurately, robustly and collaboratively to achieve natural and precise answer generation in the end.
4.3.1. Question Decomposition
The question decomposition stage processes the user question and identifies specific query triggers for the sensor data query stage, such as the context of interest, date and time ranges, and the summarization function, as illustrated in the leftmost box in Fig. 4. The primary challenge in designing this stage lies in generating accurate decompositions for arbitrary user inputs, including various questions types as discussed in Sec. 3. We choose to rely on LLMs for this stage due to their outstanding capability in handling natural language inputs. However, LLMs are not without limitations as they can produce erroneous outputs or hallucinations (Huang et al., 2023), especially in our case where a dedicated dataset for similar decomposition tasks is unavailable.
To address this challenge, we propose a few-shot learning approach to prompt pre-trained LLMs, enhanced with in-context learning (Alayrac et al., 2022; Shao et al., 2023) and chain of thought techniques (Chu et al., 2024; Lu et al., 2022). An example prompt is illustrated in Fig. 6(a), where LLM is instructed to mark different arguments with distinct symbols, such as ”¡¡” and ”¿¿” for function names. This helps in accurately extracting various arguments from the LLM output. In SensorChat, we accommodate a variety of real-life scenarios by extracting contexts (such as activities), dates, times of day, and summarization functions during the query process, as shown in Fig. 6(a). SensorChat supports multiple extracted phrases to handle complex questions such as ”How long did I work at school on Monday and Tuesday?” SensorChat is designed to be flexible, allowing for future expansion with additional decomposition terms.
We next explain more details on the in-context learning and chain-of-thought techniques used in SensorChat, both for ensuring high-quality question decompositions.
In-Context Learning (ICL) integrates a few examples directly into the prompt during inference, enabling LLMs to adapt effectively to specific tasks without requiring fine-tuning (Alayrac et al., 2022; Shao et al., 2023). Building on this insight, we design a library of general solution templates covering diverse QA scenarios and incorporate them as in-context examples to improve decomposition accuracy (Fig. 6(a)). We observe that decomposition solutions for the same question type often share similarities. For instance, ”how long” questions typically map to the CalculateDuration function. To utilize this question-specific property, in SensorChat, we first classify the question type using a BERT model (Devlin et al., 2018). Depending on the question type, we dynamically select templates that best match the question category, a strategy shown to enhance ICL performance (Fu et al., 2023). Fig. 6(b) illustrates an example of a solution template for a time query. Including an explanation in the solution template is particularly helpful for complex reasoning tasks.
Chain-of-Thought (CoT) Prompting explicitly requests LLMs to generate their reasoning process, which improves accuracy in zero- or few-shot logical reasoning (Chu et al., 2024; Lu et al., 2022). In SensorChat, we include “please generate step-by-step explanations” in prompts to encourage CoT, as shown in Fig. 6(a). CoT and detailed explanations in solution templates improve the quality of logical reasoning, especially for questions requiring multi-step thoughts. For example, applying CoT to the question “What did I do right after waking up on Wednesday?” helps produce a step-by-step thinking process, such as: “To find the activity right after waking up on Wednesday, we can query the start and end times of all activities on Wednesday, therefore we can use the function ¡¡DetectFirstTime¿¿ and ¡¡DetectLastTime¿¿ with ((all activities)) on [[Wednesday]]”.
(a) The prompt design for the question decomposition stage includes ICL and CoT.
(b) Example solution template for the time query question category.
(c) Prompt design in the answer assembly stage for LLM finetuning.
Figure 6. Prompts design in SensorChat to better leverage the power of LLMs.
4.3.2. Sensor Data Query
The sensor query stage is the core module of SensorChat, responsible for searching the sensor database and extracting relevant information, which directly impacts the accuracy of SensorChat’s answers. Existing sensor-based QA systems did not employ an explicit search module, restricting to low-dimensional sensor features (Englhardt et al., 2024; Kim et al., 2024; Yang et al., 2024b) or fixed windows of sensor signals (Xing et al., 2021; Moon et al., 2023b, a; Chen et al., 2024; Arakawa et al., 2024). In contrast, SensorChat introduces this explicit query stage to handle long-duration, fine-grained timeseries sensor signals and provide accurate information to queries, which is not possible with prior QA designs.
The sensor and label encoders are pretrained offline as explained in Sec. 4.2. With the pretrained sensor encoder, the full-history raw sensor signals are converted into an embedding database beforehand. During online user interactions, as shown in the middle part of Fig. 4, the sensor query stage only needs to encode the context (e.g., activity) generated in the question decomposition stage using the text encoder (④ in Fig. 4) , performs a similarity search among the sensor database given the text (⑤), and summarizes the retrieved sensor embeddings into textual information for the final answer assembly (⑥). The design of the sensor query stage faces two primary challenges: (1) ensuring accurate and efficient searches within the sensor embedding database, and (2) constructing adequate output context based on users’ queries. SensorChat addresses these challenges through an efficient similarity-based embedding search mechanism and a set of carefully designed summarization functions, as detailed below.
Similarity-based Embedding Search Our goal is to accurately identify all sensor samples relevant to the query context. For example, if the query context is “exercise”, SensorChat aims to retrieve all sensor samples associated with activities similar to exercise. If multiple text are generated during question decomposition, SensorChat performs an embedding search for each query text individually. The pretraining in Sec. 4.2 ensures that the sensor embedding space is well aligned with the text embedding space even using partial context. Therefore, the similarity comparison given any pair of sensor and text embeddings can be achieved by training a similarity function $f$ in the embedding space:
(2) $f(\mathbf{z}^{s}_{t},\mathbf{z}^{w})=f\big{(}\theta(\mathbf{x}_{t}),\phi(w)\big{)}\in[0,1]$
The output of $f$ is a scalar value between 0 and 1, representing the similarity between the sensor and text samples. With $f$, the similarity search in the embedding space is significantly more efficient than searching directly through raw, high-dimensional time-series data. The efficiency can be further improved by narrowing the search scope based on the date and time-of-day arguments identified during question decomposition.
Summarization Function Design We design a set of summarization functions to ”summarize” the queried sensor samples and generate contextual information for the answer assembly stage. The specific summarization function to be used is determined by SensorChat during question decomposition. For example, a question of ”how long” should be directed to the CalculateDuration function, while a question of ”what did I do” should be handled by the DetectActivity function. Each summarization function uses a unique template to return text information. The numerical value in the returned text is determined by the embedding search results. For instance, when querying CalculateDuration with the activity “cooking”, date “Sunday”, and time “morning”, the text output can be: ”You spent $\gamma$ minutes cooking on Sunday morning”, where $\gamma$ is calculated as follows:
(3) $\gamma=\sum_{t\in T_{SundayMorning}}\Bigg{[}f\big{(}\theta(\mathbf{x}_{t}),\phi(\textrm{"cooking"})\big{)}>h\Bigg{]}.$
Here $T_{SundayMorning}$ represents all timestamps within Sunday morning. The notation $[Cond]$ gives $1$, when the inner condition Cond is met; otherwise 00. $h$ is a predetermined threshold.
In SensorChat, we carefully design a set of summarization functions to account for diverse scenarios in real life, including time quries, activity quries, counting, etc. The details of those functions are explained in Appendix A.
4.3.3. Answer Assembly
As shown in the right box of Fig. 4, the final stage of answer assembly integrates question and sensor information to generate the final answer. State-of-the-art methods (Xing et al., 2021; Zhang et al., 2023; Moon et al., 2023b) rely on “black-box” fusion of natural language and sensory data, often leading to ineffective fusion and inaccurate answers (see Sec. 3). In contrast, SensorChat summarizes query results to text and directly fuses them with the question in the prompt, as illustrated in Fig. 6(c). Our intuition is that, in contrast to processing and fusing with other modalities, LLMs are the most professional in dealing with text. SensorChat is capable of answering both qualitative and quantitative questions by combining the original question and the extracted fine-grained activity information from long-duration, high-dimensional sensors. At this answer assembly stage, we finetune a LLM such as LLaMA (Zhang et al., 2023) to adapt the model to the desired answer style. Fine-tuning is chosen over few-shot learning as it delivers better performance with the presence of high-quality datasets like SensorQA (Reichman et al., 2025) (see Sec.6.6). We use Low Rank Adaptation (LoRA) (Hu et al., 2021) due to its parameter efficiency and comparable performance to full fine-tuning.
5. System Implementation
We implement SensorChat on real-world systems. We envision SensorChat as a personal assistant that provides accurate and timely answers to user questions, as outlined in our problem statement (Sec. 3.1). Fig. 7 visualizes the general system pipeline of SensorChat. We employ smartphones and smartwatches to collect multimodal sensor data from users in daily lives (① in Fig. 7(a)). Implemented based on the ExtraSensory App (Vaizman et al., 2018a), the mobile devices automatically gathers data for 20 seconds every minute, including 40Hz IMU signals, 13 MFCC audio features from a 22kHz microphone, and other phone state information including compass, GPS location, Wi-Fi status, light intensity, battery level, etc. The details can be found in the ExtraSensory App manual (Vaizman et al., 2018a). These sensor data are then transmitted to a system running SensorChat (② in Fig. 7(a)). We offer two variants of SensorChat, designed for a cloud server and an edge environment respectively. Their detailed implementations and trade-offs are discussed below. Finally, users can interact with SensorChat directly through a chatting interface using natural language, shown as ③ in Fig. 7(a).
SensorChatC and SensorChatE We offer two system variants of SensorChat, as shown in Fig. 7(b) and 7(c).
  • SensorChatC, designed for a cloud environment, uses GPT-3.5-Turbo (Ye et al., 2023) for question decomposition and a full-size finetuned LLaMA2-7B model (Touvron et al., 2023a) for answer assembly. We deploy and test SensorChat on a cloud server equipped with an NVIDIA A100 GPU (a10, 2025).
  • SensorChatE, designed for an edge environment, uses quantized LLaMA model (Touvron et al., 2023a) for both question decomposition and answer assembly. The question decomposition model is quantized from the official LLaMA3-8B, while the answer assembly model is quantized from our fine-tuned version of LLaMA2-7B. We use Activation-aware Weight Quantization (AWQ), a state-of-the-art quantization method for LLMs, known for its hardware efficiency. We deploy and test SensorChatE on a NVIDIA Jetson Orin NX module (jet, 2025) with 16GB RAM.
SensorChatC and SensorChatE accommodate two typical use scenarios. SensorChatC is expected to deliver superior QA performance with the full-precision LLMs in the cloud, at the cost of intensive resource consumption. Additionally, SensorChatC requires the users to transmit the full sensor history to the cloud server. On the other hand, SensorChatE runs entirely on a local edge platform belonging to the user, eliminating the need to transmit user data to the cloud and thus preserving user privacy. However, its QA and latency performance degrade compared to SensorChatC.
(a) Real system implementation.
(b) SensorChatC diagram.
(c) SensorChatE diagram.
Figure 7. System implementation details of SensorChat.
Implementation Details The algorithm part of SensorChat is implemented with Python and PyTorch (Paszke et al., 2019). For the offline sensor encoder, SensorChat uses a Transformer architecture (Vaswani, 2017) with 6 encoder layers, 8 attention heads, and a feedforward network size of 2048 for time series data. For the low-dimensional phone status data, SensorChat employs a fully connected layer as the encoder. The fusion layer is also implemented as a fully connected layer. The label encoder is initialized from the pretrained CLIP ViT-B/32 label encoder (Radford et al., 2021). Both the sensor and label embedding spaces share a dimension size of 512. Offline pretraining is performed with the partial-context loss proposed in Sec. 4.2 using a temperature scalar of $\tau=0.1$. We use the Adam optimizer with a learning rate of $51e-$ 5 over 100 epochs.
For question decomposition, we design two solution templates for each question category. Limiting the number of templates to two helps balance prompt length and performance. For online data queries, we initialize the classifier $f$ as a multilayer perceptron with one hidden layer and a ReLU activation function. The hidden layer size is set to 512 and the query threshold is configured to $h=0.5$. For finetuning the LLM in answer assembly, we apply low rank adaptation (LoRA) finetuning (Hu et al., 2021) on an A100 GPU, using a batch size of 8, a learning rate of 0.0002 and 10 epochs. The LoRA rank $r$ is 16, the LoRA scaling factor alpha is 16, and the LoRA dropout is 0.1. During answer generation, SensorChat uses a max sequence length of 1024 and a geberating temperature of $0.2$.
6. Evaluation on State-of-the-Art Dataset
In this section, we thoroughly evaluate SensorChat on the state-of-the-art dataset focusing on quantitative questions. We further validate SensorChat in a real world study with open-ended, qualitative questions in Sec. 7.
6.1. Dataset and Metrics
In this evaluation section, we focus mainly on the SensorQA dataset (Reichman et al., 2025) and quantitative questions. A detailed introduction to SensorQA (Reichman et al., 2025) is provided in Sec. 3. To the best of our knowledge, SensorQA is the first and only available benchmarking dataset for QA interactions that use long-term timeseries sensor data and reflect practical user interests. While we focus on SensorQA (Reichman et al., 2025) in this section, we emphasize that the motivation and design of SensorChat are broadly applicable and can be extended to other practical sensing applications. To ensure the best alignment between the QA pairs and sensor information, we conduct offline encoders pretraining on the ExtraSensory multimodal sensor dataset (Vaizman et al., 2017), which servers as the sensor data source for SensorQA. During pretraining, all sensor samples are aligned by a time window of 20 seconds.
Dataset Variants and Metrics We evaluate three versions of SensorQA (Reichman et al., 2025) using various metrics to assess both the quality and accuracy of the generated answers.
  • Full answers refer to the original full responses in SensorQA. We evaluate the model’s performance on the full answers dataset using Rouge-1, Rouge-2, and Rouge-L scores (Eyal et al., [n. d.]). Rouge scores measure the overlap of n-grams between the machine-generated content and the ground-truth answers, expressed as F-1 scores. Higher Rouge scores indicate greater similarity between the generated and true answers.
  • Short answers are the 1-2 key words extracted from the full answers by GPT-3.5-Turbo (Ye et al., 2023), offered with the original SensorQA dataset (Reichman et al., 2025). We use the exact match accuracy on the short answers to evaluate the precision of generated answers, as detailed in Sec. 3.
  • Multiple choices are generated by prompting GPT-3.5-Turbo (Ye et al., 2023) to create three additional choices similar to the correct short answer. An example QA can be ”Which day did I spend the most time with coworkers? A. Friday, B. Monday, C. Thursday, D. Wednesday”, with the correct answer being ”D” or ”D. Wednesday.” The models are expected to accurately select the correct answer from the four candidates. We evaluate the performance based on exact answer selection accuracy.
We create the multiple-choice version in addition to the full answers and short answers provided in the original SensorQA dataset (Reichman et al., 2025), to further assess the model’s ability in distinguishing similar facts based on sensor data. We use different dataset variants to assess various aspects of the models. The full answers dataset evaluates overall language quality, while the short answers and multiple-choice evaluations focus on the model’s ability to learn underlying facts rather than relying solely on patterns in language token generation. Further discussion of the evaluation metrics can be found in Sec. 8.
6.2. State-of-the-Art Baselines
We compare SensorChat against several state-of-the-art baselines that leverage different modalities combinations, including text, vision+text, and sensor+text data. These comparisons highlight the effectiveness of integrating multiple modalities to address the sensor-based QA tasks. We consider both closed-source and open-source baselines for a comprehensive analysis.
We first evaluate the state-of-the-art closed-source generative models using various modalities:
  • GPT-3.5-Turbo (Ye et al., 2023) and GPT-4 (Achiam et al., 2023) are text-only baselines taking only the questions as inputs.
  • GPT-4-Turbo (Achiam et al., 2023) and GPT-4o (Achiam et al., 2023) are vision+text baselines taking the activity graphs and the questions as inputs. For all vision+text baselines, we feed the activity graphs in SensorQA (Reichman et al., 2025), similar to Fig. 2, along with the questions into the model.
  • IMU2CLIP+GPT-4222https://github.com/facebookresearch/imu2clip (Moon et al., 2023b) is the state-of-the-art sensor+text GPT baseline as explained in Sec. 3.
For these closed-source generative models, we use few-shot learning (FSL). Specifically, we incorporate a set of QA examples from SensorQA (Reichman et al., 2025) into the prompt for each question input. We adopt $10$ samples per question based on a grid search of $\{2,5,10,15\}$.
We further conduct comprehensive evaluation with the state-of-the-art open-source models using various modalities:
  • T5 (Raffel et al., 2020) and LLaMA [^3] https://www.llama.com/ (Touvron et al., 2023a) are popular text-only language models.
  • LLaMA-Adapter [^4] https://github.com/OpenGVLab/LLaMA-Adapter (Zhang et al., 2023) is a recent vision+text framework offering a lightweight method for fine-tuning instruction-following and multimodal LLaMA models. It integrates vision inputs (i.e., activity graphs from SensorQA (Reichman et al., 2025)) with LLMs using a transformer adapter. We utilize the latest LLaMA-Adapter V2 model.
  • LLaVA-1.5555https://llava-vl.github.io/ (Liu et al., 2024a) represents state-of-the-art vision+text model. LLaVA connects pre-trained CLIP ViT-L/14 visual encoder (Radford et al., 2021) and large language model Vicuna (Chiang et al., 2023), using a projection matrix. LLaVA-1.5 (Liu et al., 2024a) achieves state-of-the-art performance on 11 benchmarks through simple modifications to the original LLaVA and the use of more extensive public datasets for finetuning.
  • DeepSQA [^6] https://github.com/nesl/DeepSQA (Xing et al., 2021) trains a CNN-LSTM model with compositional attention to fuse sensor+text modalities and predict from a fixed and limited set of candidate answers given questions and IMU signals. We adapt their implementation to use the full-history timeseries data as input, to align with SensorChat’s setup.
  • OneLLM [^7] https://github.com/csuhan/OneLLM (Han et al., 2023) is a state-of-the-art multimodal LLM framework that processes sensor+text modalities using a universal pretrained CLIP encoder and a mixture of projection experts for modality alignment. We adapt their implementation and feed the full-history timeseries data from the IMU tokenizer.
All baselines based on LLaMA (that is, LLaMA, LLaMA-Adapter, and OneLLM) use LLaMA2-7B (Touvron et al., 2023a). For T5 and DeepSQA, we train the models directly on the SensorQA dataset (Reichman et al., 2025). For the LLM baselines, we apply LoRA fine-tuning (FT) (Hu et al., 2021) using the samples from SensorQA (Reichman et al., 2025). For all methods that require training, we randomly select 80% of the QA pairs in SensorQA (Reichman et al., 2025) as training samples and reserve the remaining 20% for testing. We explore alternative splitting schemes in Sec. 6.8 to demonstrate SensorChat’s generalizability to unseen users. All baselines adopt the same hyperparameters as those specified in their official codebases. We did not compare with the latest works of Sensor2Text (Chen et al., 2024), PrISM-Q&A (Arakawa et al., 2024), and DrHouse (Yang et al., 2024b) due to limited access to their open-source code and models.
(a) Example of a time query question on a single-day duration.
(b) Example of a counting question on a multi-day duration.
Figure 8. Qualitative results of SensorChat in comparison to state-of-the-art methods.
6.3. Qualitative Performance
We illustrate a qualitative comparison of SensorChat with the top-performing baselines in Fig. 8. Specifically, Fig. 8 (a) focuses on time-related queries and (b) focuses on counting questions, which are the most challenging for SOTA methods given the long-duration time series sensor inputs (see Sec. 3). Key phrases in the answers are highlighted in the green if they closely match the true answer, and in the red if they do not. The two presented examples are very challenging for state-of-the-art baselines. SensorChat, on the other hand, consistently produces more accurate answers which can be attributed to its novel three-stage pipeline.
Answering the question in Fig. 8 (a) requires two major steps: (i) calculating the total time spent in school and in the main workplace, and (ii) computing the difference between them. SensorChat accurately answers it by first decomposing and then querying the durations for “at school” and “in the main workplace” respectively, resulting in “You spent 11 hours and 27 minutes at school Tuesday. You spent 9 hours and 3 minutes at main workplace”. Finally, SensorChat integrates the above text and determines the time difference during the answering stage. Although SensorChat’s answer is 15 minutes off from the true value, possibly due to inaccuracies in the sensor encoder or LLM reasoning, SensorChat still approximates the ground truth with an accuracy unmatched by other baselines.
In Fig. 8 (b), counting the total days spent at home requires long-term reasoning where LLaMA-Adapter (Zhang et al., 2023) and DeepSQA (Xing et al., 2021) usually fall short. SensorChat decomposes this question and queries the duration of ”at home” on ”each day”, then leaving the counting task to the answering stage. In a nutshell, SensorChat relies on question decomposition and sensor data queries to extract relevant key sensor information, while the answer assembly stage handles reasoning and produces the final answer. This collaboration across the three stages allows SensorChat to effectively manage a wide range of tasks, particularly those requiring multi-step reasoning and quantitative analyzes, which highlights SensorChat’s advancements over existing works.
Modalities Method FSL/FT1 Full Answers Short Answers Multiple Choices
Rouge-1 ($\uparrow$) Rouge-2 ($\uparrow$) Rouge-L ($\uparrow$) Accuracy ($\uparrow$) Accuracy ($\uparrow$)
Text GPT-3.5-Turbo (Ye et al., 2023) FSL 0.35 0.23 0.32 0.03 0.33
Text GPT-4 (Achiam et al., 2023) FSL 0.66 0.51 0.64 0.16 0.34
Text T5-Base (Raffel et al., 2020) FT 0.71 0.55 0.69 0.25 0.52
Text LLaMA2-7B (Touvron et al., 2023b) FT 0.72 0.62 0.72 0.26 0.56
Vision+Text GPT-4-Turbo (Achiam et al., 2023) FSL 0.38 0.28 0.36 0.14 0.24
Vision+Text GPT-4o (Achiam et al., 2023) FSL 0.39 0.28 0.37 0.20 0.07
Vision+Text LLaMA-Adapter (Zhang et al., 2023) FT 0.73 $0.57$ $0.71$ 0.28 0.54
Vision+Text LLaVA-1.5 (Liu et al., 2024a) FT $0.62$ $0.46$ $0.60$ 0.21 0.47
Sensor+Text IMU2CLIP-GPT4 (Moon et al., 2023b) FSL 0.44 0.28 0.40 0.13 0.18
Sensor+Text DeepSQA (Xing et al., 2021) FT 0.34 0.05 0.34 0.27 -
Sensor+Text OneLLM (Han et al., 2024) FT 0.12 0.04 0.12 0.05 0.30
Sensor+Text SensorChatC FT 0.77 0.62 0.75 0.54 0.70
Sensor+Text SensorChatE FT 0.76 0.60 0.74 0.49 0.67
1FS: Few-Shot Learning. FT: Finetuning.
Table 4. Quantitative results of SensorChat compared against state-of-the-art methods. Bold and underlined values show the best results (all achieved by SensorChat) and the best among baselines.
6.4. Quantitative Performance
Table 4 presents the quantitative results of all methods on the three variants of SensorQA (Reichman et al., 2025): full answers, short answers, and multiple-choice. Note that DeepSQA (Xing et al., 2021) was not evaluated on the multiple-choice version due to its inability to handle dynamic answer choices. It is important to highlight that exact match accuracy for short and multiple-choice answers is a strict metric, as it requires the model to generate answers in the exact same form as the correct ones. For instance, ”4 hours” and ”3 hours 50 min” would be considered different. Answering multiple-choice questions can also be particularly challenging, as candidate answers differ only slightly, such as ”A. 10 min” vs. ”B. 20 min,” making them difficult to distinguish. In practical applications, however, a QA system does not necessarily need to achieve perfect exact match accuracy to be useful. We leave the exploration of more advanced metrics that better align with user satisfaction for future work, as discussed in Sec. 8.
As shown in Table 4, our SensorChat outperforms the best state-of-the-art methods with the highest Rouge scores on full answers, 26% higher accuracy on short answers, and 14% higher accuracy on multiple choices. The top Rouge scores highlight SensorChat’s ability to generate high-quality natural language that are the most similar with the ground-truth answers in SensorQA (Reichman et al., 2025). Even under the strict exact match evaluation, SensorChat achieves 54% accuracy on short answers and 70% on multiple-choice questions. These accuracy improvements demonstrate SensorChat’s effectiveness in learning the underlying activity facts from the long-duration, multimodal time series sensor data. As explained in Sec. 6.3, all three stages are crucial for achieving high accuracy, including correct question decomposition to invoke the right functions, precise sensor queries to accurately extract activity information from sensor data, and effective answer assembly for generating natural language responses. SensorChatC performs slightly better than SensorChatE mainly due to the stronger model capability of GPTs compared to quantized LLaMA in question decomposition. However, SensorChatE, as a pure edge solution, preserves better user privacy as discussed in Sec. 5.
In contrast, all baselines struggle with quantitative accuracy, with an highest accuracy of merely 28% on the short answers. Despite GPT-4’s strong reasoning capabilities and its multimodal variants (GPT-4-Turbo and GPT-4o), the models do not perform optimally for processing sensor data and daily life activities, as they are not specifically trained for these tasks. The Rouge scores are lower because GPTs tend to generate longer text, resulting in less similarity with the ground-truth answers. Interestingly, text-only baselines with finetuning, such as LLaMA2-7B (Touvron et al., 2023a) and T5 (Raffel et al., 2020), achieve some of the highest accuracy among the baselines, with 27% accuracy on short answers and 56% accuracy on multiple-choice datasets. These results suggest a sensor bias in the dataset, where some questions can be “guessed” correctly without sensor data. Accuracy scores lower than these baselines indicate ineffective fusion of text and sensor data, as seen with models like Llava-1.5 (Liu et al., 2024a) and OneLLM (Han et al., 2024), whose poor performance stems from mismatches in pretraining and finetuning data formats. For example, OneLLM was designed for aligning raw IMU signals with fixed window sizes, making it challenging to adapt to long-duration sensor data in our task. Finally, DeepSQA (Xing et al., 2021) and LLaMA-Adapter (Zhang et al., 2023) perform best among the baselines but struggle with accurate quantitative analysis on long-duration time series, as discussed in Sec.3 and Sec.6.3.
Figure 9. The answer accuracy and latency trade-offs of all methods measured on the cloud and edge platforms. We evaluate SensorChatC on A100 (a10, 2025) and SensorChatE on Jetson Orin NX (jet, 2025). All LLM-based models are quantized to 4-bit weights with AWQ (Lin et al., 2024a) for Jetson Orin deployments.
6.5. End-to-End Answer Generation Latency
To evaluate efficiency, we measure the end-to-end answer generation latency of SensorChatC and SensorChatE on their respective platforms - cloud server and NVIDIA Jetson Orin. For a fair comparison, we quantize all LLM-based baselines to 4-bit weights using AWQ (Lin et al., 2024a) for the Jetson Orin experiments, matching SensorChatE ’s setup. SensorChat takes an average generation latency of 2.3s on the cloud and 10.0s on the Jetson Orin. Specifically, SensorChatC takes an average of 1.2s for question decomposition, 0.5s for sensor data query, and 0.6s for answer assembly. SensorChatE takes an average of 2.5s, 4.9s, and 2.5s for each stage, respectively.
The accuracy-latency trade-off of SensorChat and locally running baselines is shown in Fig. 9. Despite higher latency due to its dual-LLM design, SensorChat outperforms other baselines in accuracy. SensorChat still achieves real-time responses on the cloud and can be deployed on resource-constrained devices with reasonable generation latency. Notably, quantization causes negligible accuracy loss in SensorChatE compared to the full-precision SensorChatC. This is due to SensorChat’s design, which converts sensor data into text, making the final answer assembly a purely language-based task. Techniques like AWQ (Lin et al., 2024a) minimize accuracy degradation in language tasks after quantization. We plan to further optimize the latency of SensorChat on edge devices via algorithm-system co-design in future work.
6.6. Ablation Studies
In this section, we comprehensively examine the impact of key design choices in SensorChat. Without loss of generality, we use SensorChatC as the base model.
Table 5. Impact of three major stage in SensorChatC. Bold values highlight the best results.
Setup Full Answers Short Answers
Rouge-1 ($\uparrow$) Rouge-2 ($\uparrow$) Rouge-L ($\uparrow$) Accuracy ($\uparrow$)
w/o Question Decomposition 0.73 0.57 0.71 0.35
w/o Sensor Data Query 0.72 0.62 0.72 0.26
w/o Answer Assembly 0.26 0.08 0.24 0.0
Full SensorChat 0.77 0.62 0.75 0.54
Impact of Each Stage in SensorChat We first evaluate the individual contribution of each stage in SensorChat by removing one of them from the pipeline. By removing question decomposition, we use a fixed and general decomposition templates for all questions. Removing sensor data query reverts the model to a language-only approach. By removing answer assembly, we directly output the queried sensor context from the second stage. Table 5 summarizes the results on full and short answers including both quality and quantitative accuracy. As observed, removing any stage leads to a significant performance drop. In SensorChat, all three stages must work collaboratively to deliver high-quality, accurate answers across diverse question types in SensorQA. Among the three stages, the answer assembly stage has the most significant impact, as it directly influences the final output. Removing it results in severely degraded performance, with near-zero accuracy on short answers. However, the question decomposition and sensor data query stages are equally crucial.
Table 6. Impact of various designs for sensor-text pretraining in SensorChatC. Bold values highlight the best results.
Sensor Data Training Loss Online Querying Multiple Choices
Accuracy ($\uparrow$) Answer Accuracy ($\uparrow$)
Statistical features Partial-Context Loss 0.91 0.62
Time series IMU2CLIP (Moon et al., 2023b) 0.90 0.61
Time series Partial-Context Loss 0.98 0.70
Impact of Sensor Features and Pretraining Loss Functions We next evaluate the impact of sensor features and loss functions during offline encoder pretraining. Specifically, we compare using statistical features (e.g., mean acceleration) versus raw time series inputs, and IMU2CLIP loss (Moon et al., 2023b) versus our proposed contrastive sensor-text pretraining loss for partial context (see Sec.4.2). For IMU2CLIP, text samples are generated by combining all valid labels into one sentence. We report the online querying accuracy and multiple-choice answer accuracy to assess the influence to sensor information extraction. As shown in Table 6, statistical features result in low querying and answer accuracies. This validates our motivation to design SensorChat that high-dimensional time series sensor data are critical for fine-grained activity information. IMU2CLIP training, even with fine-grained data, yields poorer querying and answering accuracies, highlighting its limited ability associating sensor embeddings from partial text query. Our proposed loss function, which aligns sensor and text encoders for partial context queries, proves more effective. These findings emphasize the importance of selecting appropriate sensor features and loss functions during pretraining in order to achieve high-performance QA.
Impact of LLM Design Choices We finally evaluate the impact of various design choices for LLMs. In question decomposition, we assess the contribution of in-context learning (ICL), chain-of-thought (CoT) techniques, and different backbone LLMs. The results are summarized in Table 7. Both ICL and CoT are crucial for high-quality and accurate answers. This is because an effective question decomposition improves sensor data queries. The solution templates in ICL are more essential to SensorChat as removing ICL reduces answer accuracy by 11%. CoT enhances reasoning and slightly boosts accuracy by 1-5%. Interestingly, using a more advanced backbone (GPT-4 vs. GPT-3.5-Turbo) results in minimal improvement in answer quality, as GPT-4, while generating richer text, does not follow instructions as well as GPT-3.5-Turbo according to our observation.
For answer assembly, we evaluate the effectiveness of finetuning compared to zero-shot or few-shot learning, as well as different LLaMA backbones. As shown in Table 7, zero-shot learning results in poor performance, while few-shot learning improves answer quality but still lags behind finetuning. This highlights that finetuning is the most effective approach when a dataset like SensorQA is available. Using a more advanced LLaMA backbone, such as LLaMA3-8B, has minimal impact. Finetuning and the dataset prove to be more important than the model architecture during answer assembly.
Table 7. Impact of various design choices for LLMs in SensorChatC. Bold values highlight the best results.
Stage Setup Model in Stage Full Answers Short Answer
Rouge-1 ($\uparrow$) Rouge-2 ($\uparrow$) Rouge-L ($\uparrow$) Accuracy ($\uparrow$)
w/o ICL GPT-3.5-Turbo 0.75 0.59 0.72 0.43
Question w/o CoT GPT-3.5-Turbo 0.76 0.61 0.74 0.50
Decomposition Full GPT-4 0.77 0.62 0.75 0.49
Full GPT-3.5-Turbo 0.77 0.62 0.75 0.54
Zero-shot learning LLaMA2-7B 0.16 0.07 0.14 0.0
Answer Few-shot learning LLaMA2-7B 0.43 0.29 0.41 0.24
Assembly Finetuning LLaMA3-8B 0.76 0.61 0.74 0.53
Finetuning LLaMA2-7B 0.77 0.62 0.75 0.54
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10. Sensitivity of key hyperparameters.
6.7. Sensitivity Analyses
Fig. 10 shows the sensitivity of key parameters in SensorChat. while the less sensitive ones are omitted due to space limitation. The default parameter setting is the same as described in Sec. 5. We mainly focus on evaluating the short answers to assess the parameters’ impact on factual information extraction.
Learning Rate in Answer Assembly Fig. 10 (leftmost) shows the short answers accuracy for learning rates of $\{1e-5,5e-5,1e-4,2e-4,5e-4\}$ during finetuning. Larger learning rates result in bigger gradient steps during LoRA finetuning, with $41e-$ 4 providing the best performance for our task.
LoRA Rank in Answer Assembly Fig. 10 (middle left) shows the short answers accuracy for LoRA ranks of $\{8,16,32,64\}$ during finetuning. The rank affects the size of the LoRA adapter weights. Higher ranks mean more parameters and a larger weight space to optimize. For our task, varying ranks have little impact on final accuracy, with rank 32 achieving the best performance for short answers.
Generating Temperature in Answer Assembly Fig. 10 (middle right) shows the short answers accuracy for generating temperatures of $\{0.01,0.1,0.2,0.5\}$ during the final answer generation. Higher temperatures instruct the LLM to use more “creativity”. For our task, varying temperatures have negligible impact on the short answers accuracy, indicating minimal impact to the sensor information extraction.
Query Threshold in Sensor Data Query Fig. 10 (rightmost) shows the F1 scores of online querying at different query thresholds $h=\{0.2,0.4,0.5,0.6,0.8\}$. Raising the threshold $h$ excludes less confident positive predictions, which can improve F1 scores. However, this may also overlook some detailed events, potentially reducing answer accuracy. Ideally, $h$ should be calibrated individually for each user to achieve the best results.
(a) Average short-answer accuracy on different users.
(b) Average online querying accuracy on different users during sensor data query.
Figure 11. Generalization of key learning components in SensorChat.
6.8. Generalization and Robustness
We evaluate SensorChat’s generalization and robustness across different users’ sensor data and QA inputs. In addition to the standard 80/20 random split, we compare results with a split where training is performed on the first 48 out of 60 users and testing includes all users. To ensure a fair comparison, we equalize the training set size in both splits by duplicating samples in the smaller set. Our evaluation focuses on two key learning processes in SensorChat: LLM fine-tuning in answer assembly and sensor-text encoder pretraining.
Fig. 11(a) presents the short-answer accuracy when fine-tuning on all users’ QA data versus only the first 48 users. The results show similar accuracy for both seen and unseen users, demonstrating SensorChat’s strong generalizability in answer assembly. This is likely due to SensorChat’s design of treating answer assembly as a pure language task. Since all users’ language tokens follow similar distributions in a sensor-based QA task, generalization remains robust across user variations.
Fig.11(b) presents the online querying accuracy when pretraining with all users’ sensor data versus only the first 48 users. Unlike language fine-tuning, limiting sensor data to the first 48 users leads to accuracy degradation on unseen users due to variations in data distributions. Therefore, improving SensorChat’s generalization to new users primarily depends on developing a robust sensor and text encoder, which is a well-studied problem in existing literature (Xu et al., 2023b). We leave the investigation for combining with these techniques in future work.
7. Real User Study
We conduct a user study to evaluate SensorChat’s real-world applicability. In addition to dataset-based evaluations in Sec. 6 which focus on quantitative questions, this study places SensorChat into the “wilderness”. Participants will interact with SensorChat immediately after data collection and may pose arbitrary questions, especially the qualitative and open-ended ones. Our study is approved by the Institutional Review Board Committee.
User Study Setup We recruited eight volunteers (five males and three females) who were instructed to follow their normal daily routines while carrying a smartphone with the ExtraSensoryApp (Vaizman et al., 2018a). The smartphone models include Huawei Mate 10 Pro, LG G7 ThinQ and Google Pixel 2. The app automatically collected multimodal sensor data and transmitted it to SensorChat whenever a network connection was available. Reporting activity labels was optional for the participants. The data collection phase lasted one to three days, with valid samples ranging from 52 to 1366 minutes. Sensor data availability varied due to factors such as phone model and usage patterns. After the data collection phase, volunteers were invited to interact with SensorChatC in-person through a chat-based graphical interface. They were encouraged to ask any questions about their lives during the data collection phase and observe SensorChatC’s response generation in real time. Finally, we gathered feedback from the participants including ratings on answer content, latency, and practical value of SensorChat on a scale from 1 to 5. We specifically select SensorChatC for the user study to evaluate SensorChat’s full functionality in real-world conditions. We leave user evaluation and optimization of SensorChatE in a purely edge scenario for future work.
User Feedback Results Fig. 12 displays the quantitative feedback ratings gathered from the eight participants. SensorChat received an average score of 3.12 for answer content, 4.50 for latency, and 4.00 for practical utility, supporting SensorChat’s applicability in real life scenarios. Participant comments praised SensorChat’s natural responses, e.g., “The answers were formatted in an easy to understand way”. However, concerns were also raised about quantitative accuracy, such as “some numbers were a little off” and “it mentioned activities I never did”. We attribute these issues to the limited performance of the sensor encoder when generalizing from a dataset to real-world users, as we observed noisy outputs from the sensor query stage, occasionally detecting activities that never occurred. Fortunately, generalizing models trained on one dataset to different settings is a well-studied problem (Xu et al., 2023b). Integrating these techniques into SensorChat could enhance its practical performance, which we leave for future investigation.
In terms of latency, all participants give positive feedbacks regarding the end-to-end answer generation latency, with comments such as ”I do not feel as if I had to wait a long time for the answers” and ”I think it is faster than I thought previously”. Most participants are positive in terms of the practical utility of SensorChat in their daily lives, e.g., “I believe it can help with my wellness management significantly.” The less positive comments mentioned the challenge of adapting SensorChat to individual users for generating more personalized and useful responses, an issue that we leave for future exploration.
Generalization to Qualitative and Open-Ended Questions Participants have asked creative and open-ended questions to SensorChat that do not exist in SensorQA, providing a more comprehensive assessment of SensorChat’s practical utility. For example, one participant asked, “How would you rate my lifestyle and what improvements do you suggest?” SensorChat responded with, “I would rate this lifestyle as 2 out of 5 stars and would suggest you improve on your exercise and talking,” based on the duration of each activity where exercise and talking were lacking. The three-stage pipeline, especially the explicit sensor data query stage, enables SensorChat to analyze users’ overall lifestyle and fitness, in addition to answering quantitative questions in SensorQA (Reichman et al., 2025). Specifically, SensorChat achieves this through comprehensively querying a list of activity schedule in the sensor data query stage and then analyzing them with LLM during the answer assembly stage. Participants appreciated SensorChat’s ability to effectively handle both quantitative and qualitative questions, e.g., ”I think it could be very useful to be able to answer these qualitative and quantitative questions about my lifestyle,” highlighting its broad applicability to diverse real-life scenarios.
Figure 12. Satisfactory ratings from eight participants on three questions.
8. Discussion & Future Work
Generalizability of SensorChat In the quantitative study, we mainly focus on SensorQA (Reichman et al., 2025) as it is the only available sensor-based QA dataset that aligns with our problem statement. However, SensorChat is not limited to this dataset. As demonstrated in the user study (Sec.7), SensorChat can generalize to broader real-life scenarios and answer high-level qualitative questions. Moreover, the design of SensorChat can be easily expanded. Specifically, SensorChat can incorporate more sensors via adding additional modality encoders. It can also adapt to more diverse questions by expanding the solution templates or using more powerful LLMs such as OpenAI’s o3-mini (ope, 2025) or DeepSeek (dee, 2025) during question decomposition. Last but not least, SensorChat can serve a larger user base by collecting more data and equipping with better generalization designs such as (Xu et al., 2023b). SensorChat is a general and flexible framework designed for fusing long-duration multimodal sensors knowledge with language and answering versatile questions.
Personalization of SensorChat This work emphasizes SensorChat’s ability to generalize across a broad range of questions and users. However, providing personalized responses is equally crucial, particularly for health-related analyses. We foresee the potential to tailor SensorChat for individual users by personalizing the LLMs (Kim et al., 2024), a direction we intend to explore in future research.
Evaluation Metrics SensorChat uses traditional NLP metrics and accuracy to evaluate the quality and precision of generated answers. While these metrics excel in objectiveness, they do not capture subjective user satisfaction, which is crucial for assessing a system’s readiness for a broad market. We leave a comprehensive evaluation of SensorChat with subjective metrics as our future work.
Efficiency of SensorChat In this work, we demonstrate that SensorChat can be deployed on Jetson-level edge devices using common quantization techniques. However, further optimization is needed for real-time LLM service on edge devices. The long-term goal for systems like SensorChat is to run on mobile devices. Recent studies (Liu et al., 2024b; Zhuang et al., 2024) have explored enabling real-time LLM serving on mobile devices, and these techniques could be integrated into SensorChat in future work. Additionally, we recognize the potential synergy between SensorChat and vector databases (Zhou et al., 2024) or retrieval-augmented generation (Zhao et al., 2024). The techniques in these domains can be integrated into SensorChat to improve query performance and efficiency.
9. Conclusion
Natural interactions between users and multimodal sensors can unlock the full potential of sensor data in real-world applications. However, existing systems struggle to process long-duration, high-dimensional sensor data, often resulting in unsatisfactory answers to long-term questions. In this paper, we present SensorChat, the first edge system designed to handle long-duration, time-series sensor data for both qualitative and quantitative question answering. SensorChat introduces a novel three-stage pipeline including question decomposition, sensor data query and answer assembly. All stages and interfaces are carefully designed to effectively fuse the natural language and the sensor modality. Evaluation results and user studies demonstrate SensorChat’s effectiveness in answering a wide range of qualitative and quantitative questions. The cloud version of SensorChat system supports real-time interactions, while the edge version is optimized to run on Jetson-level devices.
Acknowledgements.
This work was supported in part by National Science Foundation under Grants #2003279, #1826967, #2100237, #2112167, #1911095, #2112665, #2120019, #2211386 and in part by PRISM and CoCoSys, centers in JUMP 2.0, an SRC program sponsored by DARPA.
Appendix A Summarization Function Design
We design the following summarization functions to handle various real-life QA scenarios.
  • CalculateDuration is used for time comparisons, time queries, and existence questions.
  • DetectActivity handles activity queries. CountingFrequency and CountingDays handle single- and multi-day counting.
  • DetectFirstTime and DetectLastTime are used for specific timestamp queries.
We emphasize that the function set can be expanded in the future to accommodate a broader range of queries. However, since the summarization function is selected by the LLM during question decomposition, it is crucial to limit the number of functions to avoid overwhelming or confusing the model. Fortunately, one strength of SensorChat is its ability to transform questions into existing functions during question decomposition, eliminating the need to create new ones unnecessarily. For instance, a question like ”What did I do right after waking up?” can be decomposed into calls to DetectFirstTime and DetectLastTime, rather than requiring the creation of a new function such as DetectNextActivity. To enable this transformation, we include a list of available functions in the prompt, as shown in Fig. 6(a). This approach ensures that the current set of six summarization functions is sufficient to effectively handle the vast majority of user queries.